I have a linear equation problem that I don't understand like y=2x-5 to find the y coordinates of the points (-2),(4), and (-1) How do I solve this one step by step.



I have another one like 2x + y = -6 how do I solve this one too. Step by step

Answer by sofiyac(983)

so it's giving you x you need to find y
so y=2x-5
first point is x=-2 so plug it in
y=2*(-2)-5
y=-4-5
y=-9
so first coordinate is (-2,-9)
you do the same thing with the rest, so next one is x=4
y=2*4-5
y=8-5
y=3
(4,3) and so on